OpenStudy (ashley1nonly):

) Carbon-14 has a half-life of 5720 years and this is a first order reaction. If a piece of wood has converted 75% of the carbon-14, then how old is it? A) 11440 years B) 2375 years C) 4750 years D) 4290 years E) 1430 years I NEED THE STEPS TO SOLE FOR A TEST IN TWO WEEKS

OpenStudy (cuanchi):

do you know what a half life means?

OpenStudy (ashley1nonly):

the time it takes for the thing to half its life

OpenStudy (cuanchi):

to half its concentration!!

OpenStudy (ashley1nonly):

which is what the [A] means, the concentration

OpenStudy (cuanchi):

The problem said that has lost already the 75% how much it is left?

OpenStudy (ashley1nonly):

5720*.75=4290

OpenStudy (cuanchi):

no!!!

OpenStudy (ashley1nonly):

5720-4290=1430

OpenStudy (cuanchi):

in the first half life will lost 50% of the original amount

OpenStudy (cuanchi):

in the second half life is going to lost the 50% of the 50%

OpenStudy (ashley1nonly):

So we are supposed to have the 5720, then take 25%

OpenStudy (cuanchi):

so that means that in the second half life will lost an additional 25%

OpenStudy (ashley1nonly):

5720*.5 =2860*.25=715 2860-715= 2145

OpenStudy (cuanchi):

no!!!

OpenStudy (ashley1nonly):

how would you work it out

OpenStudy (cuanchi):

you lost already 75%, that's means you have only 25% of the original

OpenStudy (cuanchi):

after the first half life (5720 years) you lost 50%

OpenStudy (cuanchi):

in a second half life (5720 years) you will lost 25% more of the original

OpenStudy (cuanchi):

so after 2 half life you will have only 25% of the original C-14

OpenStudy (ashley1nonly):

understood and that should leave us with 11440 which is the answer???

OpenStudy (ashley1nonly):

we would'nt us ln[A]=-kt+ln[A]0 that's whats involved with the first order

OpenStudy (cuanchi):

that is correct, this is a easy way to do it, the other way when the % is not 50% (1 half life), 25% (2 half lifes, 12.5% (3 half lifes), 6.25% (4 half lifes) you have to use the t half and the linearized equation for a first order reaction to calculate the time

OpenStudy (ashley1nonly):

so we are solving for t

OpenStudy (cuanchi):

yes first you calculate the k with the half life equation and then you replace the [A]o -f for the percentage 100 and whichever other

OpenStudy (cuanchi):

yes and solve for t

OpenStudy (ashley1nonly):

[A}0 would be 75 ln[A] would be the 5720 and k would be ...

OpenStudy (cuanchi):

in fact the k = ln2/ half life

OpenStudy (cuanchi):

no!!!

OpenStudy (cuanchi):

[A]o= 100 [A]f= 25 you lost 75% !!!

OpenStudy (ashley1nonly):

ok since we are at 75% we start at 5720, then we double at 50%. so wouldnt it just be 5720*2 since we passed two half lives which would give it its age

OpenStudy (cuanchi):

we are not at 75%, we are a 25%

OpenStudy (cuanchi):

11440 years ago was 100%

OpenStudy (cuanchi):

and in 5720 years from now will still have 12.5% of the C-14

OpenStudy (ashley1nonly):

I see what you mean now. [A]0 is what we started with whivh was 100 [A]f is what we were left which (100-75=25) so [A]f is 25. then by redoing the equation we can say ln[A]f/[A]0=-kt

OpenStudy (cuanchi):

yes!!!,

OpenStudy (ashley1nonly):

Woot woot thanks sooo much.... i know i was diffuclt